OpenStudy (coplops):

HELP ME, RIP.How many solutions does this linear system have? y =2/3 x+ 2 6x – 4y = –10 one solution: (–0.6, –1.6) one solution: (–0.6, 1.6) no solution infinite number of solutions

OpenStudy (raffle_snaffle):

What do you think?

OpenStudy (coplops):

Don't I need to solve y= 2/3x+2

OpenStudy (raffle_snaffle):

you need to solve for x or y. Take that equation and plug into other equation to solve for either x or y.

OpenStudy (raffle_snaffle):

y = 2/3x + 2 6*x - 4(2/3x + 2) = -1 solve for x. Then take that value sand plug into equatin to solve for y.

OpenStudy (raffle_snaffle):

not -1, suppose to be -10.

OpenStudy (raffle_snaffle):

Then you will have your solution.

OpenStudy (coplops):

6*x - 4(2/3x + 2) = -1 and when I get that answer I plug it into y

OpenStudy (raffle_snaffle):

No, you are solving for x, you plug the x value into one of the equations to solve for y.

OpenStudy (raffle_snaffle):

y = 2/3x + 2 doesn't need to be solved. I mean you could solve for x, but you are just doing more work than needed. This equation is solved for y. Sub. the equation here in for the y into the other equation. When doing this we have one equation in terms of x. Now you can solve for x.

OpenStudy (raffle_snaffle):